### v2

@@ 1 Introduction
Topological entropy
of a shift of finite type, which in statistical physics is usually called free energy per site, is the asymptotic growth rate of the number of restrictions on $\llbracket 0,n-1\rrbracket^{d}$ of its elements. In a celebrated article, E.H.Lieb [Lie67] computed an exact formula of topological entropy for the square ice model, with the rationale of developing tools for computing efficiently some physically relevant quantities for models with high number of variables. Unfortunately, the method proposed by Lieb does not generalize to other shifts of finite type easily. In general, computing exactly topological entropy of a multidimensional shift of finite type is a very hard problem.
@@
Together with M. Sablik [GS21], the first author proposed a quantification of the block gluing property, in which a function of the size of the patterns, called gap function, replaces the constant. This function reflects the ‘range of order’ in the system: the larger this function, the farther the presence of one particular pattern has an influence over the configuration. They studied the ‘edge of uncomputability’ (by analogy with the edge of chaos) for this quantification, with the purpose of understanding how the uncomputability phenomenon appears. In particular, they identified the ‘area’ in between logarithmic and linear functions as where uncomputability occurs for topological entropy. Unfortunately, no tool is available in order to analyze shifts of finite type in this area. In particular, it is not known if there exists a shift of finite type in two dimensions which has the block gluing property with a gap function strictly between logarithmic and linear. On the class of decidable shifts, close to the one of shifts of finite type, the first and second authors identified [GdM19] a threshold at which uncomputability of entropy occurs, defined by the condition
@@
In order to understand better block gluing classes, which group together shifts of finite type having equivalent gap functions, the strategy that we propose here is to restrict the scope to a natural subclass of the one of two-dimensional shifts of finite type. In this text, we consider Hom shifts, that is, the set of graph morphisms from $\mathbb{Z}^{d}$ to $G$, where $G$ is an undirected, simple, connected graph (self-loops are allowed). In the symbolic dynamics context, they were studied by N. Chandgotia [Cha17], who coined the name ‘Hom shifts’, borrowing the definition from G. Brightwell and P. Winkler [BW00]. Although entropy is computable on Hom shifts [Fri97], this is a natural class to better understand the block-gluing property for the following reasons. 1. First, it is related to statistical physics models, in the sense that the hard square shift is included in it, and the square ice model is related to the set of three-colourings of $\mathbb{Z}^{2}$ (in particular they have the same entropy), also in this class. 2. While several problems are undecidable for multidimensional shifts of finite type, many of them become decidable for Hom shifts. For instance, Hom shifts are defined by a symmetric set of forbidden patterns, and such subshifts have algorithmically computable entropy [Fri97] - a fact related to the intuition that it is impossible to embed universal computation in shifts under this (strong) constraint. 3. The conceptual richness of graph theory should help to forge concepts in order to analyze block gluing classes in this restricted context, concepts which may then be extended to the general context of multidimensional shifts of finite type.
@@
In particular, we infirm R. Pavlov and M. Schraudner’s conjecture. In order to prove this theorem, we extend the methods developed in [CM18], and remove the square free hypothesis by considering, instead of the universal cover, its quotient by squares of $G$.
@@ 2 Definitions and notations
For any set $S$, we denote by $S^{*}$ the set of finite words on $S$. For a word $u$, we denote the number of its letters by $|u|$. We usually write $u$ as $u_{0}\ldots u_{|u|-1}$. The empty word is denoted by $\epsilon$. Let us denote *=$\mathbb{N}^{*}=\{1,2,\dots\}$ the set of positive integers and set =*$\mathbb{N}=\mathbb{N}^{*}\cup\{0\}$. For all integers $a,b\in\mathbb{Z}$, we denote by $\llbracket a,b\rrbracket$ the interval $\{j\in\mathbb{Z}:a\leq j\leq b\}$. Similarly $\rrbracket a,b\llbracket:=\llbracket a+1,b-1\rrbracket$. Let also $\|\cdot\|_{\infty}$ be the norm defined by $\|\textbf{k}\|_{\infty}:=\max(|\textbf{k}_{1}|,|\textbf{k}_{2}|)$ for all $k\in\mathbb{Z}^{2}$.
@@ 2.2 Graphs
A walk on the graph $G$ is a non-empty word $p$ in $V_{G}^{*}$ such that for all $k\leq|p|-2$, $(p_{k},p_{k+1})\in E_{G}$. We denote by $l(p)$ the number $|p|-1$, and call it the length of $p$ (equivalently, this is the number of edges that the walk follows). A cycle on $G$ is a walk $c$ such that $c_{0}=c_{l(c)}$. It is said to be simple when $i<j$ and $c_{i}=c_{j}$ imply that $i=0$ and $j=l(c)$. Similarly we say that a walk $p$ is simple when for all $i\neq j$, $p_{i}\neq p_{j}$.
@@
We denote by $\varphi$ the function from the set of walks on $G$ to itself defined as follows: for every walk $p$, $\varphi(p)$ is obtained from $p$ by replacing successively every spine $aba$ by $a$ until there is none left. $\varphi$ is well-defined because the remaining word does not depend on the order in which the spines are replaced.
@@ 2.3 Hom shifts
For every walk $p$ and right or left shift $q$ of $p$, we have $d_{G}(p,q)=1$.
@@ 3.1 Distance between walks
Because $\mathbb{U}$ is block-like, a point $(i,j)\notin\mathbb{U}$ cannot have more than two neighbors in $\mathbb{U}$. Let us assume that there is some $(i,j)\notin\mathbb{U}$ which has two neighbors in $\mathbb{U}$. These neighbors are respectively of the form $(i\pm 1,j)$ and $(i,j\pm 1)$, otherwise $\mathbb{U}\cap(\{j\}\times\mathbb{Z})$ and $\mathbb{U}\cap(\mathbb{Z}\times\{j\})$ would not be an interval. Without loss of generality, let us assume that they are $(i+1,j)$ and $(i,j+1)$ (the other cases are dealt with similarly). Let us prove that $(i+1,j+1)\in\mathbb{U}$. Since $\mathbb{U}$ is connected, there exists a walk on $\mathbb{U}$ from $(i+1,j)$ to $(i,j+1)$. Such a walk intersects $\{i\}\times\rrbracket-\infty,j\llbracket$ or $\{i+1\}\times\rrbracket j,+\infty\llbracket$ (see for instance Figure 1).
@@
Every block-like subset of $\mathbb{Z}^{2}$ such that no $(i,j)\notin\mathbb{U}$ has two neighbors in $\mathbb{U}$ is a rectangle. Indeed, for every $j\in\mathbb{Z}$ such that the columns ×$\mathbb{Z}\times\{j\}$ and ×$\mathbb{Z}\times\{j+1\}$ intersect $\mathbb{U}$ non-trivially, their respective intersections with $\mathbb{U}$ are equal. A similar statement is satisfied for rows. This implies that $\mathbb{U}$ is a rectangle.
@@
[…]
@@
If $\Delta_{G}^{n}$ is not bipartite, it has a cycle of odd length. Along this cycle, if we take the last letter of the walk on $G$ corresponding to each vertex, we get a cycle in $G$ of odd length. Reciprocally, assume $G$ has a cycle $c$ of odd length $m$. For all $j$ between 00 and $m-1$, we denote by $p^{(j)}$ the walk on $G$ which begins at $c_{j}$ and alternates between $c_{j}$ and $c_{j+1}$. Then $p^{(0)}\ldots p^{(m-1)}p^{(0)}$ is a cycle of odd length on $\Delta_{G}^{n}$. ∎
@@
[…]
@@
[…]
@@
$w|_{\textbf{v}+\llbracket 0,n-1\rrbracket^{2}}=q$
@@ 3.2 From walks to cycles
The inequality $\gamma_{G}(n)\leq\gamma_{G}(n+1)$ is trivial. Let us prove the second one. Let us consider two walks $p$ and $q$ of length $n+1$. There exists $m\leq\gamma_{G}(n)$ and a walk $p^{(0)},\dots,p^{(m)}$ in $\Delta^{n+1}_{G}$ from $p^{(0)}=p_{0}\ldots p_{n+1}$ and $p^{(m)}=q_{0}\ldots q_{n+1}$. For all $i\geq 1$, let us set $q^{(i)}=p^{(i)}p^{(i-1)}_{n+1}$ and $q^{(0)}=p$. For all $i<m$, $(q^{(i)},q^{(i+1)})$ is an edge of $\Delta^{n+1}_{G}$, and $q^{(0)}=p$. Since $q^{(m)}$ is equal to $q$ except for the last vertex, $q^{(m)}\in\rho_{r}(\rho_{l}(q))$. As a consequence $d_{G}(p,q)\leq l+2$, and thus $\gamma_{G}(n+1)\leq\gamma_{G}(n)+2$. ∎
@@
[…]
@@
    There exists $m\leq\text{diam}(G)+1$ such that there is a walk from $p_{n}$ to $p_{m}$ of length $m$ or $m-1$: For all $k\in\{0,\dots,n\}$, let us set $l_{k}:=k-\delta(p_{n},p_{k})$. We have that $l_{0}=-\delta(p_{n},p_{0})\leq 0$, and $l_{\text{diam}(G)+1}\geq 1$ by definition of the diameter. Furthermore, for all $k<n$, we have:
    $0\leq l_{k+1}-l_{k}\leq 2$
    As a consequence there exists an integer $m\leq\text{diam}(G)+1$ such that $l_{m}\in\{0,1\}$, which means that $\delta(p_{n},p_{m})$ is $m$ or $m-1$.
@@
    We have $\delta(p_{n},p_{m})=m$: Let $q$ be a walk of length $\delta(p_{n},p_{m})$ from $p_{n}$ to $p_{m}$. If $q$ is of length $m-1$, the walk $p_{m}\ldots p_{n}\odot q_{0}\ldots q_{m-1}$ is a cycle of odd length $n-1$, which is impossible since $G$ is bipartite; it follows that $q$ is of length $m$.
@@ 4.1 Universal cover
In the following sections, we introduce concepts which are built upon the universal cover, in particular what we call the square cover (Section 4.3). Let us take a moment here to explain the intuition behind this construction. We would like to establish a correspondence between $X_{G}$ and $X_{\mathcal{U}_{G}}$, in order to deduce the block gluing gap function of $X_{G}$ from the one of $X_{\mathcal{U}_{G}}$: the latter is much easier since $\mathcal{U}_{G}$ is a tree.
@@
$a$$b$$a$$c$$a$$b$$a$$b$$c$$a$$c$$a$$c$$b$$a$
Figure 4: On the left: a graph $G$ containing a cycle of length four ($aabca$). On the right: the cycle $w=acaacbaacbaba$ obtained by following a loop in $\mathbb{Z}^{2}$ in a configuration of $X_{G}$.
@@ 4.2 Square decomposition
It is sufficient to see that for $p,q$ two walks which differ by a square, $l(p)-l(q)$ is even. Therefore every cycle of a square-decomposable graph is of even length. ∎
@@ 4.3 Square cover
For all $\textbf{i}\in\mathbb{Z}^{2}$, take any walk $\textbf{i}_{0}\dots\textbf{i}_{m}$ in $\mathbb{Z}^{2}$ such that $_{m}=\textbf{i}$ i , and define $z_{\boldsymbol{i}}=\pi_{x_{\boldsymbol{i}_{0}}}(x_{\textbf{i}_{0}}\dots x_{\textbf{i}_{m}})\in\mathcal{U}^{\square}_{G}[x_{\boldsymbol{i}_{0}}]$. Taking another walk in $\mathbb{Z}^{2}$ would yield the same equivalence class $z_{\boldsymbol{i}}$, since the corresponding walks in $G$ are equivalent for $\sim_{\square}$. Therefore $z$ is well-defined and satisfies the requirements. ∎
@@
Let us prove this by induction on the area of the cycle $p\odot q^{-1}$. When this area is equal to 00, $p$ is equal to $q$ and they are therefore equivalent for $\sim_{\square}$. Let us assume that the statement is proved whenever $p\odot q^{-1}$ has area $\leq m$, and fix $p,q$ for which this cycle has area $m+1$. Consider some square $s$ and integer $k$ such that $\varphi((p\odot q^{-1})\oplus_{k}s)$ has area $m$. Let us denote by $p^{\prime}$ and $q^{\prime}$ the following paths:
@@
$\eta^{-1}_{v}(w)_{i}\text{ is the class for }\sim_{\square}\text{ of the walk }\varphi(w_{1}\dots w_{i}).$
@@
Let us fix some vertex $a$ of $G$. Since $\mathcal{U}^{\square}_{G}$ is infinite and the degree of every vertex is less than $|V_{G}|$, for all $n$ there exists a path $p_{n}$ on $G$ such that $\pi_{a}(p_{n})$ is at distance $2n$ from $\pi_{a}(a)$ in $\mathcal{U}^{\square}_{G}[a]$. Let $a_{0},\ldots,a_{2n}$ be walks on $G$ such that $\pi_{a}(a_{0})\cdots\pi_{a}(a_{2n})$ is a walk (of length $2n$) from $\pi_{a}(a)$ to $\pi_{a}(p_{n})$ in $\mathcal{U}^{\square}_{G}[a]$. We denote $u:=aa_{2}\cdots a_{2n}$ its image for $\eta$, and $v:=(aa_{2})^{n}a$. Let us prove that $d_{G}(u,v)\geq n$.

### v3

@@ 1 Introduction
Topological entropy.
The topological entropy of a shift of finite type, which in statistical physics is usually called free energy per site, is the asymptotic growth rate of the number of restrictions on $\llbracket 0,n-1\rrbracket^{d}$ of its elements. In a celebrated article, E.H.Lieb [Lie67] computed an exact formula of topological entropy for the square ice model, with the rationale of developing tools for computing efficiently some physically relevant quantities for models with high number of variables. Unfortunately, the method proposed by Lieb does not generalize to other shifts of finite type easily. In general, computing exactly topological entropy of a multidimensional shift of finite type is a very hard problem.
@@
Together with M. Sablik [GS21], the first author proposed a quantification of the block gluing property, in which a function of the size of the patterns, called gap function, replaces the constant. This function reflects the ‘range of order’ in the system: the larger this function, the farther the presence of one particular pattern has an influence over the configuration. They studied the ‘edge of uncomputability’ (by analogy with the edge of chaos) for this quantification, with the purpose of understanding how the uncomputability phenomenon appears. In particular, they identified the ‘area’ in between logarithmic and linear functions as where uncomputability occurs for topological entropy. Unfortunately, no tool is available in order to analyze shifts of finite type in this area. In particular, it is not known if there exists a shift of finite type in two dimensions which has the block gluing property with a gap function strictly between logarithmic and linear. On the class of decidable shifts, close to the one of shifts of finite type, the first and second authors identified [GHdM19] a threshold at which uncomputability of entropy occurs, defined by the condition
@@
In order to understand better block gluing classes, which group together shifts of finite type having equivalent gap functions, the strategy that we propose here is to restrict the scope to a natural subclass of the one of two-dimensional shifts of finite type. In this text, we consider Hom shifts, that is, the set of graph morphisms from $\mathbb{Z}^{d}$ to $G$, where $G$ is an undirected, simple, connected graph (self-loops are allowed). In the symbolic dynamics context, they were studied by N. Chandgotia [Cha17], who coined the name ‘Hom shifts’, borrowing the definition from G. Brightwell and P. Winkler [BW00]. Although entropy is computable on Hom shifts [Fri97], this is a natural class to better understand the block-gluing property for the following reasons. 1. First, it is related to statistical physics models, in the sense that the hard square shift is included in it, and the square ice model is related to the set of three-colourings of $\mathbb{Z}^{2}$ (in particular they have the same entropy), also in this class. 2. While several problems are undecidable for multidimensional shifts of finite type, many of them become decidable for Hom shifts. For instance, Hom shifts are defined by a symmetric set of forbidden patterns, and such shifts have algorithmically computable entropy [Fri97] - a fact related to the intuition that it is impossible to embed universal computation in shifts under this (strong) constraint. 3. The conceptual richness of graph theory should help to forge concepts in order to analyze block gluing classes in this restricted context, concepts which may then be extended to the general context of multidimensional shifts of finite type.
@@
In particular, we disconfirm R. Pavlov and M. Schraudner’s conjecture. In order to prove this theorem, we extend the methods developed in [CM18], and remove the square free hypothesis by considering, instead of the universal cover, its quotient by squares of $G$.
@@ 2 Definitions and notations
For any set $S$, we denote by $S^{*}$ the set of finite words on $S$. For a word $u$, we denote the number of its letters by $|u|$. We usually write $u$ as $u_{0}\ldots u_{|u|-1}$. The empty word is denoted by $\epsilon$. Let us denote $\mathbb{N}^{*}=\{1,2,\dots\}$ the set of positive integers and set $\mathbb{N}=\mathbb{N}^{*}\cup\{0\}$. For all integers $a,b\in\mathbb{Z}$, we denote by $\llbracket a,b\rrbracket$ the interval $\{j\in\mathbb{Z}:a\leq j\leq b\}$. Similarly $\rrbracket a,b\llbracket:=\llbracket a+1,b-1\rrbracket$. Let also $\|\cdot\|_{\infty}$ be the norm defined by $\|\textbf{k}\|_{\infty}:=\max(|\textbf{k}_{1}|,|\textbf{k}_{2}|)$ for all $k\in\mathbb{Z}^{2}$.
@@ 2.2 Graphs
A walk on the graph $G$ is a non-empty word $p$ in $V_{G}^{*}$ such that for all $k\leq|p|-2$, $(p_{k},p_{k+1})\in E_{G}$. We denote by $l(p)$ the number |p|1|p|-1, and call it the length of $p$ (equivalently, this is the number of edges that the walk follows). A cycle on $G$ is a walk $c$ such that $c_{0}=c_{l(c)}$. It is said to be simple when $i<j$ and $c_{i}=c_{j}$ imply that $i=0$ and $j=l(c)$. Similarly we say that a walk $p$ is simple when for all $i\neq j$, $p_{i}\neq p_{j}$.
@@
We denote by $\varphi$ the function from the set of walks on $G$ to itself defined as follows: for every walk $p$, $\varphi(p)$ is obtained from $p$ by replacing successively every spine abaaba by $a$ until there is none left. $\varphi$ is well-defined because the remaining word does not depend on the order in which the spines are replaced.
@@ 2.3 Hom shifts
For every walk $p$ and right or left shift $q$ of $p$ with $p\neq q$, we have $d_{G}(p,q)=1$.
@@ 3.1 Distance between walks
Because $\mathbb{U}$ is block-like, a point $(i,j)\notin\mathbb{U}$ cannot have more than two neighbors in $\mathbb{U}$. Let us assume that there is some $(i,j)\notin\mathbb{U}$ which has two neighbors in $\mathbb{U}$. These neighbors are respectively of the form $(i\pm 1,j)$ and $(i,j\pm 1)$, otherwise $\mathbb{U}\cap(\{j\}\times\mathbb{Z})$ or $\mathbb{U}\cap(\mathbb{Z}\times\{j\})$ would not be an interval. Without loss of generality, let us assume that they are $(i+1,j)$ and $(i,j+1)$ (the other cases are dealt with similarly). Let us prove that $(i+1,j+1)\in\mathbb{U}$. Since $\mathbb{U}$ is connected, there exists a walk on $\mathbb{U}$ from $(i+1,j)$ to $(i,j+1)$. Such a walk intersects $\{i\}\times\rrbracket{-\infty},j\llbracket$ or $\{i+1\}\times\rrbracket j,+\infty\llbracket$ (see for instance Figure 1).
@@
Every block-like subset of $\mathbb{Z}^{2}$ such that no $(i,j)\notin\mathbb{U}$ has two neighbors in $\mathbb{U}$ is a rectangle. Indeed, for every $j\in\mathbb{Z}$ such that the columns $\mathbb{Z}\times\{j\}$ and $\mathbb{Z}\times\{j+1\}$ intersect $\mathbb{U}$ non-trivially, their respective intersections with $\mathbb{U}$ are equal. A similar statement is satisfied for rows. This implies that $\mathbb{U}$ is a rectangle.
@@
Let $H$ be a finite graph. For every $u,v\in V_{H}$ and $k\geq\textrm{diam}(H)$, there is a walk from $u$ to $v$ of length $k$ or $k+1$. If $H$ is not bipartite, then for all $k\geq 3\textrm{diam}(H)$, there is a walk from $u$ to $v$ of length $k$.
@@
If $\Delta_{G}^{n}$ is not bipartite, it has a cycle of odd length. Along this cycle, if we take the last letter of the walk on $G$ corresponding to each vertex, we get a cycle in $G$ of odd length. Reciprocally, assume $G$ has a cycle $c$ of odd length $m$. For all $j$ between 0 and m1m-1, we denote by $p^{(j)}$ the walk on $G$ which begins at $c_{j}$ and alternates between $c_{j}$ and $c_{j+1}$. Then $p^{(0)}\ldots p^{(m-1)}p^{(0)}$ is a cycle of odd length on $\Delta_{G}^{n}$. ∎
@@
Let $d_{G}(n):=\textrm{diam}(\Delta_{G}^{n})$. The shift $X_{G}$ is $(\Theta(d_{G}),2)$-phased-block-gluing. When $G$ is not bipartite, it is also $\Theta(d_{G})$-block-gluing.
@@
Let us first prove the second claim. We assume that $G$ is not bipartite. We show that $X_{G}$ is $O(d_{G})$-block gluing. Consider some $k\geq 3d_{G}(n)$ and let $p$ and $q$ be two locally admissible block patterns of size $n$ and $\textbf{u},\textbf{v}$ such that
@@
$w|_{\textbf{v}+\llbracket 0,n-1\rrbracket^{2}}=q\$
@@ 3.2 From walks to cycles
The inequality $\gamma_{G}(n)\leq\gamma_{G}(n+1)$ is trivial. Let us prove the second one. Let us consider two walks $p$ and $q$ of length $n+1$. There exists $m\leq\gamma_{G}(n)$ and a walk $p^{(0)},\dots,p^{(m)}$ in $\Delta^{n+1}_{G}$ from $p^{(0)}=p_{0}\ldots p_{n}$ to $p^{(m)}=q_{0}\ldots q_{n}$. For all $i\geq 1$, let us set $q^{(i)}=p^{(i)}p^{(i-1)}_{n}$ and $q^{(0)}=p$. For all $i<m$, $(q^{(i)},q^{(i+1)})$ is an edge of $\Delta^{n+1}_{G}$, and $q^{(0)}=p$. Since $q^{(m)}$ is equal to $q$ except for the last vertex, $q^{(m)}\in\rho_{r}(\rho_{l}(q))$. As a consequence $d_{G}(p,q)\leq m+2$, and thus $\gamma_{G}(n+1)\leq\gamma_{G}(n)+2$. ∎
@@
Let us assume that $G$ is bipartite. For every walk $p$ of even length, there exists a cycle $c$ such that $d_{G}(p,c)\leq\text{diam}(G)+1$.
@@
    There exists $m\leq\text{diam}(G)+1$ such that there is a walk from $p_{n}$ to $p_{m}$ of length $m$ or m1m-1: for all $k\in\{0,\dots,n\}$, let us set $l_{k}:=k-\delta(p_{n},p_{k})$. We have that $l_{0}=-\delta(p_{n},p_{0})\leq 0$, and $l_{\text{diam}(G)+1}\geq 1$ by definition of the diameter. Furthermore, for all $k<n$, we have:
    $0\leq l_{k+1}-l_{k}\leq 2.$
    As a consequence there exists an integer $m\leq\text{diam}(G)+1$ such that $l_{m}\in\{0,1\}$, which means that $\delta(p_{n},p_{m})$ is $m$ or m1m-1.
@@
    We have $\delta(p_{n},p_{m})=m$: let $q$ be a walk of length $\delta(p_{n},p_{m})$ from $p_{n}$ to $p_{m}$. If $q$ is of length m1m-1, the walk $p_{m}\ldots p_{n}\odot q_{0}\ldots q_{m-1}$ is a cycle of odd length n1n-1, which is impossible since $G$ is bipartite; it follows that $q$ is of length $m$.
@@ 4.1 Universal cover
In the following sections, we introduce concepts which are built upon the universal cover, in particular what we call the square cover (Section 4.3). Let us take a moment here to explain the intuition behind this construction. We would like to establish a correspondence between $X_{G}$ and $X_{\mathcal{U}_{G}}$, in order to deduce the block gluing gap function of $X_{G}$ from the one of $X_{\mathcal{U}_{G}}$: the latter is much easier to deal with since $\mathcal{U}_{G}$ is a tree.
@@
abc
aabbaaccaabbaabbccaaccaaccbbaa
Figure 4: On the left: a graph $G$ containing a cycle of length four (aabcaaabca). On the right: the cycle $w=acaacbaacbaba$ obtained by following a loop in $\mathbb{Z}^{2}$ in a configuration of $X_{G}$.
@@ 4.2 Square decomposition
It is sufficient to see that for $p,q$ two walks which differ by a square, l(p)l(q)l(p)-l(q) is even. Therefore every cycle of a square-decomposable graph is of even length. ∎
@@ 4.3 Square cover
For all $\textbf{i}\in\mathbb{Z}^{2}$, take any walk $\textbf{i}_{0}\dots\textbf{i}_{m}$ in $\mathbb{Z}^{2}$ such that $\textbf{i}_{m}=\textbf{i}$ , and define $z_{\boldsymbol{i}}=\pi_{x_{\boldsymbol{i}_{0}}}(x_{\textbf{i}_{0}}\dots x_{\textbf{i}_{m}})\in\mathcal{U}^{\square}_{G}[x_{\boldsymbol{i}_{0}}]$. Taking another walk in $\mathbb{Z}^{2}$ would yield the same equivalence class $z_{\boldsymbol{i}}$, since the corresponding walks in $G$ are equivalent for $\sim_{\square}$. Therefore $z$ is well-defined and satisfies the requirements. ∎
@@
Let us prove this by induction on the area of the cycle $p\odot q^{-1}$. When this area is equal to 0, $p$ is equal to $q$ and they are therefore equivalent for $\sim_{\square}$. Let us assume that the statement is proved whenever $p\odot q^{-1}$ has area $\leq m$, and fix $p,q$ for which this cycle has area $m+1$. Consider some square $s$ and integer $k$ such that $\varphi((p\odot q^{-1})\oplus_{k}s)$ has area $m$. Let us denote by $p^{\prime}$ and $q^{\prime}$ the following paths:
@@
[…]
@@
Let us fix some vertex $a$ of $G$. Since $\mathcal{U}^{\square}_{G}$ is infinite and the degree of every vertex is less than $|V_{G}|$, for all $n$ there exists a path $p_{n}$ on $G$ such that $\pi_{a}(p_{n})$ is at distance $2n$ from $\pi_{a}(a)$ in $\mathcal{U}^{\square}_{G}[a]$. Let $a=a_{0},\ldots,a_{2n}$ be walks on $G$ such that $\pi_{a}(a_{0})\cdots\pi_{a}(a_{2n})$ is a walk (of length $2n$) from $\pi_{a}(a)$ to $\pi_{a}(p_{n})$ in $\mathcal{U}^{\square}_{G}[a]$. We denote $u:=aw_{1}\cdots w_{2n}$ its image for $\eta$, and $v:=(aw_{1})^{n}a$. Let us prove that $d_{G}(u,v)\geq n$.